Bank MK expects the Euro to depreciate from its current level of $1.33/1 Euro to $1.27/ 1 Euro over the next 180 days. the bank borrowing and lending rates are as followed:

US Dollars: Lending rate 2.2% Borrowing rate: 2.4%
Euro: Lending rate: 1.6% Borrowing rate: 1.9%

A). Bank Mk wants to borrow 4 million euro in the bank market. What’s is the profit in US dollars to Bank AA if in fact the Euro does depreciate to $1.26/ 1 Euro in 180 days?

B). If Bank MG thinks the Euro will strengthen from $1.32/1 Euro to $1.34/ 1 Euro over the next 30 days and borrows US$ 3 million in the bank market, what is the profit in US$ to the Bank MG if in fact the Euro does strength to $1.34 in 30 days?
Answer #1

A) If Bank MK Borrows in Euro then he will deposit it in Us Dollars to gain profit, in order to gain he should do the following steps:

Step 1: Borrow in Euro Market of Euro 4 million, after 180 days

Amount to be paid back=4000000+ (4000000*1.9%*180/360) = Euro 4038000

Step 2: Convert it in Dollars = 4000000*1.33= $5320000

Step 3: Invest it in $ Market of $5320000

Amount to be received = 5320000+(5320000*2.2%*180/360)= $5378520

Now, Convert amount of Euro 4038000 it into $ = 4038000*1.26= $5087880

Gain or Profit in US Dollars = $5378520-$5087880 = $290640

B) If Bank MG Borrows in US Dollars then he will deposit it Euro to gain profit, in order to gain he should do the following steps:

Step 1: Borrow in Dollar Market of US$3 million, after 180 days

Amount to be paid back=3000000+ (3000000*2.4%*30/360) = $ 3006000

Step 2: Convert it in Euro = 3000000/1.32= $2272727.27

Step 3: Invest it in $ Market = $2272727.27

Amount to be received= 2272727.27+(2272727.27*1.6%*30/360)= Euro 2275757.57

Now, Convert amount of Euro 2275757.57 it into $ = 2275757.57*1.34= $3049515.14

Gain or Profit in US Dollars = $3049515.14-$3006000 = $43515.14

Note: Assume No. of days to be 360 Days.